OpenStudy (anonymous):

each question of a 5 mutiple choice exam has 3 possible answers.suppose a student picks an answer at random for each question. Find the probablity the student selects the correct answer on only the 1st question?

OpenStudy (dumbcow):

probability of picking correct answer is 1/3 so prob of picking wrong answer is 2/3 each question is an independent event to get #1 right and the rest wrong looks like this 1/3 * 2/3 * 2/3 *2/3 *2/3 =1/3 * (2/3)^4 =1/3 * 16/81 =16/243